# Foldable SuperNets: Scalable Merging of Transformers with Different Initializations and Tasks

### v1

@@ F.3 Merging text Transformers
In this series of experiments, we aimed to evaluate FS-Merge on a different modality: merging differently initialized text transformers. Following (verma2024merging), we used five bert-base-uncased models from the MultiBERTs reproductions (devlin2018bert; sellam2022the), each trained from a different random initialization and with a different data ordering. We fine-tuned each model on a distinct classification task from the GLUE dataset (wang2018glue), using six tasks: RTE, QQP, MNLI, MRPC, SST-2, and QNLI. The same pre-trained BERT was fine-tuned on QQP and MRPC, so their merge was not evaluated. It is worth noting that ViT is a pre-LN Transformer, while BERT is a post-LN Transformer (xiong2020layer).
@@
Baselines. We compared our method against two baselines: “average” (wortsman2022model), a simple weight averaging technique; and distillation (hinton2015distilling), which trains a single BERT to mimic the features of the last representation layer of the original models. We also reported the performance of the “original models”, representing the average accuracy of the models to be merged; and ensemble (ganaie2022ensemble), which averages the models’ outputs and then applies classification heads. Note that these last two are not valid merging methods as they use the original models directly.
@@
Table 22: Merging pairs of BERTs with 200 original texts from the training set of each dataset. We report the per-task accuracy on the test set.
@@
QQP, QNLI 91.01 71.63 563.2 82.16 82.79
F.4 Number of original training images - Vision Transformer
@@
F.5 Number of training data points - BERT
@@ G.1 FS-Merge ablation study
Our experiments suggest that in the ViT case, initialization is extremely important for training the Foldable SuperNet. When initialized randomly, FS-Merge does not converge, even with the addition of augmented data (and see Appendix H.2 discussing this issue). Only Foldable SuperNet initialized from the average or first model allows it to converge into a successful merged model.
@@ G.2 Distillation ablation study
Appendix H Discussion
H.1 Merging Vision Transformers: Local VS. Global perspective
@@ Appendix H Discussion
H.2 Merging Vision Transformer with randomly initialized Foldable SuperNet
In our experiments, we found that smart initialization of the Foldable SuperNet is crucial for FS-Merge. As common in NNs, we first tried to initialize the $M,U$ matrices using a random Gaussian distribution dependent on the hidden dimension (he2015delving).
@@ H.2 Merging Vision Transformer with randomly initialized Foldable SuperNet
This effectiveness is the reason the Foldable SuperNet’s $M,U$ matrices were modeled as a sum of low-rank matrices plus a concatenation of diagonal matrices (rather than just a low-rank matrix as in LoRa (hu2022lora)). The concatenation of diagonal matrices enables the initialization of the $M,U$ matrices using those successful methods (“first”, “average”).
H.3 Using inner features when merging Vision Transformers
In line with several distillation studies (wu2021one; zagoruyko2017paying; heo2019knowledge; heo2019comprehensive; park2019feed; liu2020adaptive), we tried to use the inner features of the models to be merged as a regularization for FS-Merge and for distillation. We focused on the features obtained after the MLP block or after the attention block. The attention features in the $l$ block of model $k$, created from the input $I_{\text{img}}^{k}$, can be written as $f^{k}_{l,\text{att}}(I_{\text{img}}^{k})\in\mathbb{R}^{T\times d}$. Then, the “inner loss” for the global FS-Merge, when handling two tasks $A$ and $B$, can be defined as follows:

### v2

@@ F.7 Merging text Transformers
In this series of experiments, we aimed to evaluate FS-Merge on a different modality: merging differently initialized text transformers. Following (Verma & Elbayad, 2024), we used five bert-base-uncased models from the MultiBERTs reproductions (Devlin et al., 2018; Sellam et al., 2022), each trained from a different random initialization and with a different data ordering. We fine-tuned each model on a distinct classification task from the GLUE dataset (Wang et al., 2019), using six tasks: RTE, QQP, MNLI, MRPC, SST-2, and QNLI. The same pre-trained BERT was fine-tuned on QQP and MRPC, so their merge was not evaluated. It is worth noting that ViT is a pre-LN Transformer, while BERT is a post-LN Transformer (Xiong et al., 2020). Moreover, unlike the ViT experiments where models were fine-tuned with frozen classification heads derived from CLIP embeddings, these experiments used trainable classification heads.
@@
Baselines. We compared our method against two baselines: “average” (Wortsman et al., 2022), a simple weight averaging technique; and distillation (Hinton et al., 2015), which trains a single BERT to mimic the features of the last representation layer of the original models. We also reported the performance of the “original models”, representing the average accuracy of the models to be merged; and ensemble (Ganaie et al., 2022), which averages the models’ outputs and then applies classification heads. Note that these last two are not valid merging methods as they use the original models directly.
@@
Table 29: Merging pairs of BERTs with 200 original texts from the training set of each dataset. FS-Merge was used with a low rank of 12. We report the per-task accuracy on the test set.
@@
QQP, QNLI 91.01 71.63 56.32 82.16 82.79
F.8 Number of original training images - Vision Transformer
@@
F.9 Number of training data points - BERT
@@ G.1 FS-Merge ablation study
Our experiments suggest that in the ViT case, initialization is extremely important for training the Foldable SuperNet. When initialized randomly, FS-Merge does not converge, even with the addition of augmented data (and see Appendix G.4 discussing this issue). Only Foldable SuperNet initialized from the average or first model allows it to converge into a successful merged model.
G.2 Knowledge Distillation ablation study
@@
G.3 Merging Vision Transformers: Local VS. Global perspective
@@
G.4 Merging Vision Transformer with randomly initialized Foldable SuperNet
In our experiments, we found that smart initialization of the Foldable SuperNet is crucial for FS-Merge. As common in NNs, we first tried to initialize the $M,U$ matrices using a random Gaussian distribution dependent on the hidden dimension (He et al., 2015).
@@ G.4 Merging Vision Transformer with randomly initialized Foldable SuperNet
This effectiveness is the reason the Foldable SuperNet’s $M,U$ matrices were modeled as a sum of low-rank matrices plus a concatenation of diagonal matrices (rather than just a low-rank matrix as in LoRa (Hu et al., 2022)). The concatenation of diagonal matrices enables the initialization of the $M,U$ matrices using those successful methods (“first”, “average”).
G.5 Using inner features when merging Vision Transformers
In line with several distillation studies (Wu et al., 2021a; Zagoruyko & Komodakis, 2017; Heo et al., 2019b; a; Park & Kwak, 2019; Liu et al., 2020; Wu et al., 2021b; 2023; Han et al., 2024), we tried to use the inner features of the models to be merged as a regularization for FS-Merge and for distillation. We focused on the features obtained after the MLP block or after the attention block. The attention features in the $l$ block of model $k$, created from the input $I_{\text{img}}^{k}$, can be written as $f^{k}_{l,\text{att}}(I_{\text{img}}^{k})\in\mathbb{R}^{T\times d}$. Then, the “inner loss” for the global FS-Merge, when handling two tasks $A$ and $B$, can be defined as follows:
@@ G.5 Using inner features when merging Vision Transformers
Appendix H FS-Merge expressive power
As shown in Section 3.2, Section 3.3, and Appendix F.2, FS-Merge significantly outperforms traditional merging methods, such as weight averaging and RegMean, when merging transformers trained from different initializations and tasks. In order to explain this, we present theoretical results that highlight FS-Merge’s superior expressive power compared to the baselines.
We examine the case of merging a pair of Multi-Layer Perceptron (MLP) models, $A$ and $B$. Let us consider the $l$-th layer in the MLP model $A$. The features at this layer, denoted by $f_{l}^{A}\in\mathbb{R}^{d\times N}$, can be expressed as follows:
$z_{l}^{A}=W_{l}^{A}f_{l-1}^{A}\,,\,f_{l}^{A}=\sigma(z_{l}^{A})\,.$ (8)
Here, $W_{l}^{A}\in\mathbb{R}^{d\times d}$ are the weights of the current linear layer, $d$ denotes the width, $N$ denotes the number of data points, $\sigma$ represents the ReLU activation function ($\sigma(x)=\max(0,x)$), and $z_{l}^{A}\in\mathbb{R}^{d\times N}$ are the pre-activation features. And similarly for model $B$.
Definition H.1.
A matrix $A\in\mathbb{R}^{d_{1}\times d_{2}}$ is said to have rank $r\leq d_{1},d_{2}$ if the dimension of its column space (or equivalently, its row space) is $r$. It implies that the maximum number of linearly independent columns (or rows) in $A$ is $r$.
Definition H.2.
Let $x$ be a data point sampled from the distribution $\mathcal{D}$, and let $f_{l}(x)$ denote the features of the $l$-th layer of an MLP computed from these inputs. We define the features $f_{l}$ as having a low rank of $r>0$ if $f_{l}(x)$ has a rank of at most $r$ for any $x\sim\mathcal{D}$.
We will assume low-rank features in several of the following sections, as it has been found that neural network features are often effectively low-rank Cai et al. (2021); Yu & Wu (2023). Such low-rank features imply that it is possible to use low-rank weights to achieve the same function (i.e., the input-output relation) of the model, even if the original weights are high-rank. In other words, if $f_{l}$ is a set of features with rank $r$, even if $W_{l}$ is full rank, we can find a different matrix $\bar{W}_{l}$ of rank $r$ such that
$\bar{W}_{l}f_{l}=W_{l}f_{l}\,.$ (9)
This holds because the action of a linear transformation (such as $W_{l}$) on a subspace can always be reproduced by another transformation of the same rank as the subspace.
Our main theoretical result, presented in Theorem H.3, demonstrates the significant expressive power of both full-rank and low-rank FS-Merge, surpassing that of the two popular merging methods, Git Re-Basin Ainsworth et al. (2023) and RegMean Jin et al. (2023).
Theorem H.3.
A summary of our theoretical results:
  1. 1.
    Given two MLP models, $A$ and $B$, where at each layer at least one has an invertible weight matrix, a full-rank FS-Merge can reconstruct any target MLP model of the same size from these models. The same holds for a low-rank FS-Merge, assuming that the target MLP features are low-rank.
  2. 2.
    Any merged MLP obtained using the RegMean or Git Re-Basin merging methods can also be constructed using full-rank FS-Merge, even without assuming invertible weights. Moreover, there are scenarios that FS-Merge can solve which RegMean or Git Re-Basin cannot. The same holds for low-rank FS-Merge, given the additional assumption of low-rank features.
As shown in Theorem H.3, full-rank FS-Merge has greater expressiveness compared to low-rank FS-Merge. However, we argue that in the context of merging large models with a limited amount of data, low-rank FS-Merge is preferable, as it introduces fewer learnable parameters and thus enhances generalization. This claim is supported by our experiments, which show that low-rank FS-Merge outperforms full-rank FS-Merge (Appendix G.5) and Knowledge Distillation (Appendix 3.2) when merging transformers.
Proof.
Part 1 of Theorem H.3 is proved in Appendix H.1 below.
Part 2 of Theorem H.3 follows from Proposition H.4 and Lemma H.5, which are proven in Appendix H.2 and Appendix H.3, respectively.
Proposition H.4.
Given the linear layers of two models, $A$ and $B$:
  1. 1.
    Any merged model obtained using RegMean or Git Re-Basin can also be constructed using full-rank FS-Merge.
  2. 2.
    Assuming that the features of at least one model have a low rank of $r$, then for any solution created by Git Re-Basin, low-rank FS-Merge with a rank of $r$ can create an equivalent solution.
  3. 3.
    Assuming that the features of both models have a low rank of $r$, any merged model obtained using RegMean can also be constructed using low-rank FS-Merge with a rank of $2r$.
Lemma H.5.
Consider the scenario of two linear layers that share the same backbone weight matrix, with each applying a different feature selection, permutation, and scaling matrix. Then, full-rank and low-rank FS-Merge can merge these layers into a single layer in a manner that allows them to reconstruct the original features from the merged one. In contrast, RegMean and Git Re-Basin cannot perfectly merge these two layers.
H.1 Proof of Part 1 of Theorem H.3: FS-Merge expressive power
The Foldable SuperNet for two linear layers is defined as follows:
$\tilde{f}_{l}(z_{l}^{A},\,z_{l}^{B})=U_{l}\sigma(M_{l}(z_{l}^{A}\,||\,z_{l}^{B}))\,$, (10)
where the input is a concatenation of the original pre-activation features $z_{l}^{A}\,||\,z_{l}^{B}\in\mathbb{R}^{2\cdot d\times N}$, and $\tilde{f}_{l}\in\mathbb{R}^{2\cdot d\times N}$ are the features reconstruction attempt. Intuitively, FS-Merge learns to compress and reconstruct the features of two different models, while also taking into account the activation function.
In the full rank vesion of FS-Merge, $M_{l}\in\mathbb{R}^{d\times 2\cdot d}$ and $U_{l}\in\mathbb{R}^{2\cdot d\times d}$ are the learnable parameters of the Foldable SuperNet. In the low rank version of FS-Merge, the $M_{l}$ and $U_{l}$ matrices are parametrized as a sum of a low-rank matrix and a concatenation of diagonal matrices. For example, in the case of $M_{l}$:
$M_{l}=M_{l}^{\text{diag}}+M^{1}_{l}M^{2}_{l}\,.$ (11)
When $r$ is the rank, $M^{1}_{l}\in\mathbb{R}^{d\times r}$, $M^{2}_{l}\in\mathbb{R}^{r\times 2\cdot d}$, and $M_{l}^{\text{diag}}\in\mathbb{R}^{d\times 2\cdot d}$ is a concatenation of two diagonal matrices, each with $d$ learnable parameters. A similar structure is proposed for $U$. In both version of FS-Merge, the merged layer is created as follows:
$W_{l}=M_{l}\begin{pmatrix}W_{l}^{A}&0\\ 0&W_{l}^{B}\end{pmatrix}U_{l-1}\,.$ (12)
Next, we wish to prove part 1 of Theorem H.3
Proof.
Let $l$ be an arbitrary layer in the MLPs. Given two linear weight matrices $W_{l}^{A},W_{l}^{B}\in\mathbb{R}^{d\times d}$, we will assume without loss of generality that $W_{l}^{A}$ is invertible. Let $W_{l}^{*}$ denote the target MLP weights. In the case where the input features for the weight $W_{l}^{*}$ are full rank, full-rank FS-Merge is capable of reconstructing the target weights using:
$M_{l}=\begin{pmatrix}W_{l}^{*}(W_{l}^{A})^{-1}&0\end{pmatrix}\,,\,U_{l}=\begin{pmatrix}I\\ I\end{pmatrix}\,$. (13)
As the merged weight will be
$W_{l}=\begin{pmatrix}W_{l}^{*}(W_{l}^{A})^{-1}&0\end{pmatrix}\begin{pmatrix}W_{l}^{A}&0\\ 0&W_{l}^{B}\end{pmatrix}\begin{pmatrix}I\\ I\end{pmatrix}=W_{l}^{*}\,.$ (14)
In the case where the input features for the weight $W_{l}^{*}$ have a low rank of $r$, there exists an equivalent weight $\bar{W}_{l}^{*}$ with rank $r$ that can generate the same features (i.e. $\bar{W}_{l}^{*}f_{l}=W^{*}_{l}f_{l}$). In this case, low-rank FS-Merge with rank $r$ can reconstruct these equivalent weights using a similar solution:
$M_{l}=\begin{pmatrix}\bar{W}_{l}^{*}(W_{l}^{A})^{-1}&0\end{pmatrix}\,,\,U_{l}=\begin{pmatrix}I\\ I\end{pmatrix}\,$. (15)
This solution can be achieved using a low-rank FS-Merge, as $U_{l}=\begin{pmatrix}I\\ I\end{pmatrix}$ can be reconstructed using its diagonal component $U_{l}^{\text{diag}}$. Additionally, the matrix $\begin{pmatrix}\bar{W}_{l}^{*}(W_{l}^{A})^{-1}&0\end{pmatrix}$ has rank $r$ or lower, meaning it can be reconstructed using the low-rank component $M_{l}^{1}M_{l}^{2}$.
This result happens generically, since we only require that for each layer, one of the models (either $A$ or $B$) has a full-rank weight matrix. This happens almost surely with respect to the Lebesgue measure, as almost all randomly chosen weight matrices will be full rank (i.e., singular matrices have measure zero).
H.2 Proof of Proposition H.4: FS-Merge expressiveness compared to RegMean and Git Re-Basin
Here we aim to demonstrate that full-rank FS-Merge can generate any solution produced by the popular merging methods Git Re-Basin Ainsworth et al. (2023) and RegMean Jin et al. (2023), even without assuming an invertible weight matrix. The same applies to low-rank FS-Merge, with a few additional assumptions.
Git Re-Basin Ainsworth et al. (2023) is an alignment-based merging method that aligns linear layers using a permutation matrix before averaging them. It defines the activation matching objective for the $l$-th layer as:
$P_{l}=\underset{P_{l}\in S_{d}}{\text{argmin}}\left\|f_{l}^{A}-P_{l}f_{l}^{B}\right\|_{2}^{2}\,$, (16)
Where $S_{d}$ is the set of all permutation matrices of size $d\times d$. A permutation matrix is a square binary matrix that has exactly one entry of 1 in each row and each column, with all other entries being 0. It then merges the weights as follows:
$W_{l}=\frac{1}{2}W_{l}^{A}+\frac{1}{2}P_{l}W_{l}^{B}P^{\top}_{l-1}\,.$ (17)
Intuitively, Git Re-Basin searches for a permutation matrix $P_{l}\in S_{d}$ that makes the features of the linear layers $f_{l}^{A}$ and $f_{l}^{B}$ as similar as possible (in terms of the Frobenius norm). Applying this permutation to the linear layer weights $W_{l}^{B}$ is equivalent to rearranging the neurons in the layer so that neurons generating similar features are aligned in the same positions.
RegMean Jin et al. (2023) is a strong averaging-based merging method, targeting the following objective:
$W_{l}=\underset{W_{l}}{\text{argmin}}\left\|W_{l}f_{l-1}^{A}-W_{l}^{A}f_{l-1}^{A}\right\|_{2}^{2}+\left\|W_{l}f_{l-1}^{B}-W_{l}^{B}f_{l-1}^{B}\right\|_{2}^{2}\,$, (18)
which has the next closed form solution:
$W_{l}=(f_{l-1}^{A}(f_{l-1}^{A})^{\top}+f_{l-1}^{B}(f_{l-1}^{B})^{\top})^{-1}(f_{l-1}^{A}(f_{l-1}^{A})^{\top}W_{l}^{A}+f_{l-1}^{B}(f_{l-1}^{B})^{\top}W_{l}^{B})\,.$ (19)
Next, we wish to prove Proposition H.4.
Proof.
Given that Git Re-Basin converged to a permutation matrix $P_{l}$, full-rank FS-Merge can produce the same solution using
$M_{l}=\begin{pmatrix}\frac{1}{2}I&\frac{1}{2}P_{l}\end{pmatrix}\,,\,U_{l-1}=\begin{pmatrix}I\\ P^{\top}_{l-1}\end{pmatrix}\,$. (20)
By additionally assuming that the features $f_{l}^{B}$ have a low rank of $r$, there exist projection matrices $C_{l},C^{\prime}_{l}\in\mathbb{R}^{d\times d}$, both of rank $r$, such that:
$P_{l}f_{l}^{B}=C^{\prime}_{l}P_{l}C_{l}f_{l}^{B}\,.$ (21)
And similarly for $f_{l-1}^{B}$. Thus, a low-rank FS-Merge can reconstruct the next solution:
$M_{l}=\begin{pmatrix}\frac{1}{2}I&\frac{1}{2}C^{\prime}_{l}P_{l}C_{l}\end{pmatrix}\,,\,U_{l-1}=\begin{pmatrix}I\\ (C^{\prime}_{l-1}P_{l-1}C_{l-1})^{\top}\end{pmatrix}\,$. (22)
Which is equivalent to the Git Re-Basin solution. This is a valid solution, as the identity matrices can be reconstructed using the diagonal components $U_{l-1}^{\text{diag}},M_{l}^{\text{diag}}$, while $C^{\prime}_{l}P_{l}C_{l},(C^{\prime}_{l-1}P_{l-1}C_{l-1})^{\top}$ are matrices with rank $r$ or lower.
In addition, full-rank FS-Merge can produce the same solution as RegMean using
$M_{l}=(f_{l-1}^{A}(f_{l-1}^{A})^{\top}+f_{l-1}^{B}(f_{l-1}^{B})^{\top})^{-1}\begin{pmatrix}f_{l-1}^{A}(f_{l-1}^{A})^{\top}&f_{l-1}^{B}(f_{l-1}^{B})^{\top}\end{pmatrix}\,,\,U_{l-1}=\begin{pmatrix}I\\ I\end{pmatrix}\,$. (23)
By additionally assuming that the features $f_{l-1}^{A},f_{l-1}^{B}$ have a low rank of $r$, a low-rank FS-Merge with rank $2r$ can also produce the solution in Eq. 23. This is because $U_{l-1}$ can be reconstructed using its diagonal component $U_{l-1}^{\text{diag}}$, and $M_{l}$ is a matrix with rank $2r$ or lower.
H.3 Proof of Lemma H.5: Expressive power case study
In the next section, we wish to prove Lemma H.5. We examine a simple scenario where traditional merging methods fail due to their reliance on basic merging rules. In contrast, we prove here that both full-rank and low-rank FS-Merge can handle this scenario.
In more detail, our scenario (Eq. 24) involves two linear models, $A$ and $B$, which generate features for two different tasks. We assume that both models share the same backbone weight matrix, $W^{*}\in\mathbb{R}^{d\times d}$, and that $X\in\mathbb{R}^{d\times N}$ represents the inputs. Each model then selects a subset of shared features using $V^{A/B}\in\mathbb{R}^{d\times d}$, which are diagonal matrices with 0/1 entries, each having a rank of $\frac{r}{2}$. The two matrices do not have 1s in the same positions. Next, the two models apply a permutation matrix $P^{A/B}\in\mathbb{R}^{d\times d}$ and a diagonal matrix with positive entries that scale the features, $D^{A/B}\in\mathbb{R}^{d\times d}$. The ReLU function $\sigma$ is used to generate the post-activation features $f^{A},f^{B}\in\mathbb{R}^{d\times N}$, which are later used by the classification head to make predictions.
$z^{A}=D^{A}P^{A}V^{A}W^{*}X\,,\,f^{A}=\sigma(z^{A})\,,\,\,\,\,z^{B}=D^{B}P^{B}V^{B}W^{*}X\,,\,f^{A}=\sigma(z^{B})\,.$ (24)
We will also denote
$z^{*}=(V^{A}+V^{B})W^{*}X\,,\,f^{*}=\sigma(z^{*})\,.$ (25)
Next, we prove Lemma H.5 in Appendix H.3.1, Appendix H.3.2 and Appendix H.3.3, demonstrating that low-rank FS-Merge can achieve a zero-loss solution, while Git Re-Basin and RegMean cannot.
H.3.1 FS-Merge
Low-rank FS-Merge is capable of reconstructing $f^{*}$, using
$M=\begin{pmatrix}(P^{A})^{\top}(D^{A})^{-1}&(P^{B})^{\top}(D^{B})^{-1}\end{pmatrix}\,,\,U=\begin{pmatrix}I\\ I\end{pmatrix}\,$. (26)
As the obtained features will be:
$\tilde{f}=\sigma(\tilde{z})=\sigma\left(M\begin{pmatrix}W^{A}&0\\ 0&W^{B}\end{pmatrix}UX\right)=\\ \sigma\left(\begin{pmatrix}(P^{A})^{\top}(D^{A})^{-1}&(P^{B})^{\top}(D^{B})^{-1}\end{pmatrix}\begin{pmatrix}D^{A}P^{A}V^{A}W^{*}&0\\ 0&D^{B}P^{B}V^{B}W^{*}\end{pmatrix}\begin{pmatrix}I\\ I\end{pmatrix}X\right)=\sigma(z^{*})=f^{*}\,.$ (27)
Note that Eq. 26 is a valid solution, as $M$ has a rank of at most $r$, and $U$ is a concatenation of diagonal matrices, and therefore, they can be parameterized by low-rank FS-Merge with rank of $r$.
Moreover, the low-rank Foldable SuperNet can perfectly reconstruct the features of both linear layers, achieving zero loss, using
$M=\begin{pmatrix}V^{A}(P^{A})^{\top}(D^{A})^{-1}&V^{B}(P^{B})^{\top}(D^{B})^{-1}\end{pmatrix}\,,\,U=\begin{pmatrix}D^{A}P^{A}V^{A}\\ D^{B}P^{B}V^{B}\end{pmatrix}\,$. (28)
Note that the operation of $D^{A/B}P^{A/B}V^{A/B}$ and $\sigma$ is commutative, and thus
$f^{A}=D^{A}P^{A}V^{A}f^{*}\,.$ (29)
The same holds for $f^{B}$. Using it, the Foldable SuperNet output is
$\tilde{f}(z^{A},\,z^{B})=U\sigma\left(M\begin{pmatrix}z^{A}\\ z^{B}\end{pmatrix}\right)=\begin{pmatrix}D^{A}P^{A}V^{A}\\ D^{B}P^{B}V^{B}\end{pmatrix}\sigma\left(\begin{pmatrix}V^{A}(P^{A})^{\top}(D^{A})^{-1}&V^{B}(P^{B})^{\top}(D^{B})^{-1}\end{pmatrix}\begin{pmatrix}D^{A}P^{A}V^{A}W^{*}X\\ D^{B}P^{B}V^{B}W^{*}X\end{pmatrix}\right)\\ =\begin{pmatrix}D^{A}P^{A}V^{A}\\ D^{B}P^{B}V^{B}\end{pmatrix}\sigma\left(\begin{pmatrix}V^{A}&V^{B}\end{pmatrix}\begin{pmatrix}V^{A}W^{*}X\\ V^{B}W^{*}X\end{pmatrix}\right)=\begin{pmatrix}D^{A}P^{A}V^{A}\\ D^{B}P^{B}V^{B}\end{pmatrix}\sigma\left(z^{*}\right)=\begin{pmatrix}D^{A}P^{A}V^{A}f^{*}\\ D^{B}P^{B}V^{B}f^{*}\end{pmatrix}=f^{A}\,||\,f^{B}\,.$ (30)
Moreover, this is a valid solution, as both $M$ and $U$ from Eq. 28 are matrices with a rank of at most $r$, and therefore, they can be parameterized by low-rank FS-Merge with rank of $r$.
We note that full-rank FS-Merge can generate any solution obtained by low-rank FS-Merge since it can use matrices of any rank. Therefore, full-rank FS-Merge is capable of handling these scenarios as well.
H.3.2 RegMean
Under our assumptions, RegMean’s Jin et al. (2023) objective is:
$W=\underset{W}{\text{argmin}}\left\|WX-D^{A}P^{A}V^{A}W^{*}X\right\|_{2}^{2}+\left\|WX-D^{B}P^{B}V^{B}W^{*}X\right\|_{2}^{2}\,.$ (31)
Which uses the Frobenius norm. For RegMean to achieve zero loss, it is required that $WX=D^{A}P^{A}V^{A}W^{*}X$. Thus
$W=D^{A}P^{A}V^{A}W^{*}XX^{\dagger}\,.$ (32)
Where we denote $\dagger$ as the Penrose-Moore pseudo-inverse. Similarly, to achieve zero error, we must also ensure that $W=D^{B}P^{B}V^{B}W^{*}XX^{\dagger}$. Using both results we obtain:
$D^{A}P^{A}V^{A}W^{*}XX^{\dagger}=D^{B}P^{B}V^{B}W^{*}XX^{\dagger}\,.$ (33)
Note that Eq. 33 is a set of rational equations, and therefore it can be transformed to a set of polynomial equations. These polynomials are not identically zero, due to the assumption $r>0$. Therefore, the set of solutions of these equations is a zero-measure set (under the Lebesgue measure). This means the chances of randomly selecting parameters (from some continuous distribution) that satisfy this equation are zero. Therefore, in general, RegMean cannot find parameters that satisfy this condition, so it lacks the expressive power to handle the case being studied.
H.3.3 Git Re-Basin
Under our assumptions, Git Re-Basin’s Ainsworth et al. (2023) objective is:
$P=\underset{P\in S_{d}}{\text{argmin}}\left\|D^{A}P^{A}V^{A}f^{*}-PD^{B}P^{B}V^{B}f^{*}\right\|_{2}^{2}\,.$ (34)
For RegMean to achieve zero loss, it is required that
$D^{A}P^{A}V^{A}f^{*}-PD^{B}P^{B}V^{B}f^{*}=0\,.$ (35)
Eq. 35 is a polynomial, and thus we can apply the same reasoning as in the previous section, leading to the conclusion that the probability of this equation holding in general is zero. Therefore, Git Re-Basin lacks the expressiveness needed to handle this case.
$\square$